# A Proof of the Weak Gravity Conjecture via Leading-Order Quantum Gravity Effects

### v1

@@ Abstract
The weak gravity conjecture posits that there should exist at least one state with a charge-to-mass ratio $q/M>1$, ensuring that an extremal charged black hole can still decay. In this paper, we study the corrections to the thermodynamics of Reissner-Nordström black holes within an effective field theory framework. The effective action originates from integrating out massless particles, including gravitons, at one-loop level. As a proof of the weak gravity conjecture, we show that these quantum gravity effects naturally lead to the super-extremality of charged black holes. It is important to note that our results are model-independent and have predictive capacities.
@@ I Introduction
In this regard, higher-derivative terms should be incorporated into the gravitational effective action as manifestations of quantum gravity effects Boulware:1985wk ; Cardoso:2018ptl . The quantum-corrected effective action can be formally written as
@@
Although significant achievements have been made in recent years Kats:2006xp ; Cremonini:2009ih ; Cremonini:2019wdk ; Cheung:2018cwt ; Noumi:2022ybv , a decisive conclusion has not been reached. Four-derivative terms such as $R^{2}$, $R_{\mu\nu}^{2}$ and $R_{\mu\nu\alpha\beta}^{2}$ have been extensively studied, and eq.(2) is utilized to place constraints on the combination of the coefficients of these curvature terms. There is a growing body of evidence in support of the WGC, coming from concrete models constructed from string theories or under certain assumptions. In these studies, the local $\mathcal{O}(R^{2})$ terms are often regarded as making the main contributions to the relevant corrections. However, the loop effects of massless particles including gravitons can dominate over them in the IR and provide the leading contributions Harlow:2022ich .
In this paper, we consider the loop effects of massless particles and show that these effects indeed reduce the mass of an extremal RN black hole. Remarkably, our results depend only on the low-energy particle spectrum of nature and are insensitive to the UV theories of quantum gravity. Thus, our work provides a model-independent proof of the WGC.
@@ IV Extermal RN black hole and the WGC
From the expression of $M_{ext}$, we can straightforwardly obtain our core result
@@
Interestingly, the coefficients conspire to produce a definite sign. Since $n_{s}$, $n_{f}$, $n_{v}$, and $n_{g}$ represent the numbers of particles in the spectrum, it is certain that $\mathcal{N}>0$. Consequently, quantum gravity effects indeed reduce the mass of an extremal charged black hole at a fixed charge $q$.
@@
The characteristic energy scale $\mu_{*}$ is determined by the UV theory of quantum gravity. Thus, the presence of $\ln\mu_{*}$ in eq.(14) indicates the UV sensitivity of our result. However, the model-dependent value of $\mu_{*}$ does not affect our conclusion because we are working within the regime $r>1/\mu_{*}$, where the effective field-theoretical description is valid. The physics near the scale $r\sim 1/\mu_{*}$ would be described by a complete theory of quantum gravity and is not relevant to our current analysis.
@@
$\left(\frac{\partial M_{\text{ext}}(q,\epsilon)}{\partial\epsilon}\right)_{q}\!=\!-\!\lim\limits_{M\rightarrow M_{\text{ext}}}T\left(\frac{\partial S(M,q,\epsilon)}{\partial\epsilon}\right)_{M,q},$ (16)
In fact, from the extended first law $dM=TdS+\Phi dq+U_{\epsilon}d\epsilon$, it can be easily observed that both sides of eq.(16) are equal to $\lim\limits_{T\rightarrow 0}U_{\epsilon}=\lim\limits_{T\rightarrow 0}\frac{\partial F}{\partial\epsilon}$ Xiao:2023two . Therefore, $\frac{\partial M_{ext}}{\partial\epsilon}$ can be evaluated as the spatial integral $-\int d^{3}x\sqrt{g}L_{\text{hd}}$ for an extremal RN black hole. This provides a rapid way to check whether the WGC is satisfied for a given higher-derivative term $\epsilon L_{\text{hd}}$ without delving into the analysis of the perturbed black hole thermodynamics.
@@
The WGC is mainly concerned with the fate of extremal RN black holes. However, we can also explore the quantum gravity effects on non-extremal RN black holes to see how the charge-to-mass ratio changes. In this regard, we find that
@@
The singularity at $\bar{r}_{h}=\sqrt{3}q$ indicates a critical point that coincides with the boundary between the thermodynamically stable and unstable RN black holes, as determined by the sign of the heat capacity $C_{q}$. Thus, for a thermodynamically stable black hole, the value of $q/M$ increases, while for an unstable one, it decreases.
@@ V Concluding remarks
In this paper, we have explored the effective action obtained by integrating out massless particles at the one-loop level. Our findings demonstrate that quantum gravity effects can naturally induce super-extremality ($q>M$) in charged black holes. This result offers a model-independent proof of the WGC, suggesting that quantum gravity plays a crucial role in facilitating the decay of extremal black holes into more fundamental components. These effects thus have far-reaching implications for our understanding of the ultimate evolution of black holes and the universe.
We emphasize once more that our results are independent of specific UV theories of quantum gravity. Instead, they depend solely on the low-energy particle spectrum of nature, which endows our work with predictive power. Specifically, by using the particle spectrum of the Standard Model with a minimally coupled Higgs and taking $\mu_{*}$ as the Planck energy, we predict that for a black hole with a charge of approximately $10^{-16}C$ (the typical charge carried by a dust particle), the quantity $\frac{q}{M}-1\approx 5.2\times 10^{-5}$.
In addition to proving the WGC, our study uncovers a connection between the super-extremality and the logarithmic entropy corrections of black holes. For non-extremal RN black holes, we also demonstrate a relationship between the shift in the charge-to-mass ratio and the thermodynamic stability properties of the black holes. Both of these aspects are worthy of further study.
@@ Acknowledgments
YX would like to thank Yu Tian and Hongbao Zhang for useful discussions. YX is also thankful to the Higgs Centre for Theoretical Physics at the University of Edinburgh for providing research facilities and hospitality during the visit. This work was supported in part by China Scholarship Council with Grant No. 202408130101, the Hebei Natural Science Foundation with Grant No. A2024201012, and the National Natural Science Foundation of China with Grant No. 12475048.

### v2

@@ Abstract
The weak gravity conjecture posits that there should exist at least one state with a charge-to-mass ratio $q/M>1$, ensuring that an extremal charged black hole can still decay. In this paper, we study the corrections to the thermodynamics of Reissner-Nordström black holes within an effective field theory framework. The effective action originates from integrating out massless particles, including gravitons, at one-loop level. We demonstrate that these quantum gravity effects naturally give rise to the super-extremality of charged black holes, which aligns with a recent result by Arkani-Hamed et al. Furthermore, our formalism provides a comprehensive thermodynamic analysis of both extremal and non-extremal RN black holes, offering additional interesting implications.
@@ I Introduction
In this regard, higher-derivative terms should be incorporated into the gravitational effective action as manifestations of quantum gravity effects Boulware:1985wk ; Cardoso:2018ptl ; BarrosoVarela:2023ull . The quantum-corrected effective action can be formally written as
@@
Significant achievements have been gained in recent years Kats:2006xp ; Cremonini:2009ih ; Cremonini:2019wdk ; Cheung:2018cwt ; Cheung:2019cwi ; Noumi:2022ybv . Specifically, four-derivative terms such as $R^{2}$, $R_{\mu\nu}^{2}$ and $R_{\mu\nu\alpha\beta}^{2}$ have been extensively studied, and eq.(2) is utilized to place constraints on the combination of the coefficients of these curvature terms. There is a growing body of evidence in support of the WGC, coming from concrete models constructed from string theories or under certain assumptions. In these studies, the local $\mathcal{O}(R^{2})$ terms are often regarded as making the main contributions to the relevant corrections. However, the loop effects of massless particles including gravitons can dominate over them in the IR and provide the leading contributions Harlow:2022ich .
Recently, the loop effects of massless particles have been studied by Arkani-Hamed et al. in Ref. Arkani-Hamed:2021ajd . In this paper, we will start from a non-local effective action that captures these loop effects and demonstrate that such effects indeed reduce the mass of an extremal RN black hole. This result confirms the findings of Arkani-Hamed:2021ajd , offering a fresh and concise perspective on the problem. Moreover, we present a detailed analysis of the thermodynamics of both extremal and non-extremal RN black holes. We emphasize that the results are model-independent, as they depend solely on the low-energy particle spectrum of nature and are insensitive to the UV details of quantum gravity theories.
@@ II One-Loop Effective Action
Besides, the higher curvature terms (3) shares some similarities with that analyzed in Ref.Arkani-Hamed:2021ajd through field redefinitions, and it may be equivalent in some deeper sense, as both ultimately describe the same physics of loop effects. However, the technical challenges faced in analyzing eq.(3) are different. The primary difficulty here lies in how to handle the non-local operator $\ln\frac{\Box}{\mu_{*}^{2}}$ and its variation; see, e.g., Donoghue:2014yha ; Calmet:2019eof ; Xiao:2021zly .
@@ IV Extermal RN black hole and the WGC
From the expression of $M_{ext}$, we can straightforwardly obtain the core result
@@
Interestingly, the coefficients conspire to produce a definite sign. Since $n_{s}$, $n_{f}$, $n_{v}$, and $n_{g}$ represent the numbers of particles in the spectrum, it is certain that $\mathcal{N}>0$. Consequently, quantum gravity effects indeed reduce the mass of an extremal charged black hole at a fixed charge $q$, as first pointed out in Ref. Arkani-Hamed:2021ajd .
@@
The characteristic energy scale $\mu_{*}$ is determined by the UV theory of quantum gravity. Thus, the presence of $\ln\mu_{*}$ in eq.(14) indicates the UV sensitivity of our result. However, the model-dependent value of $\mu_{*}$ does not affect our conclusion for the WGC because we are working within the regime $r>1/\mu_{*}$, where the effective field-theoretical description is valid. The physics near the scale $r\sim 1/\mu_{*}$ would be described by a complete theory of quantum gravity and is not relevant to our current analysis.
@@
$\left(\frac{\partial M_{\text{ext}}(q,\epsilon)}{\partial\epsilon}\right)_{q}\!=\!-\!\lim\limits_{M\rightarrow M_{\text{ext}}}T\left(\frac{\partial S(M,q,\epsilon)}{\partial\epsilon}\right)_{M,q}.$ (16)
In fact, from the extended first law $dM=TdS+\Phi dq+U_{\epsilon}d\epsilon$, it can be easily observed that both sides of eq.(16) are equal to $\lim\limits_{T\rightarrow 0}U_{\epsilon}=\lim\limits_{T\rightarrow 0}\frac{\partial F}{\partial\epsilon}$ Xiao:2023two . Therefore, $\frac{\partial M_{ext}}{\partial\epsilon}$ can be evaluated as the spatial integral $-\int d^{3}x\sqrt{g}L_{\text{hd}}$ for an extremal RN black hole Cheung:2018cwt ; Cheung:2019cwi . This provides a rapid way to check whether the WGC is satisfied for a given higher-derivative term $\epsilon L_{\text{hd}}$ without delving into the analysis of the perturbed black hole thermodynamics.
@@
The WGC is mainly concerned with the fate of extremal RN black holes. However, we can also explore the quantum gravity effects on non-extremal RN black holes to see how the charge-to-mass ratio changes Solomon:2020pja . In this regard, we find that
@@
The singularity at $\bar{r}_{h}=\sqrt{3}q$ indicates a critical point that coincides with the boundary between the thermodynamically stable and unstable RN black holes, as determined by the sign of the heat capacity $C_{q}$. Thus, for a thermodynamically stable black hole, the value of $q/M$ increases, while for an unstable one, it decreases. A general discussion of the correlation between thermodynamic stability and super-extremality can be found in Cheung:2018cwt , where it was first formalized based on tree-level higher derivative terms.
@@ V Concluding remarks
In this paper, we have explored the effective action obtained by integrating out massless particles at the one-loop level. We confirm that the loop effects of quantum gravity can naturally induce super-extremality ($q>M$) in charged black holes Arkani-Hamed:2021ajd . It suggests that quantum gravity plays a crucial role in facilitating the decay of extremal black holes into more fundamental components. These effects thus have far-reaching implications for our understanding of the ultimate evolution of black holes and the universe.
We emphasize that these results are independent of specific UV theories of quantum gravity. Instead, they depend solely on the low-energy particle spectrum of nature, which endows them with predictive power. Specifically, by using the particle spectrum of the Standard Model with a minimally coupled Higgs and taking $\mu_{*}$ as the Planck energy, we predict that for a black hole with a charge of approximately $10^{-16}C$ (the typical charge carried by a dust particle), the quantity $\frac{q}{M}-1\approx 5.2\times 10^{-5}$.
Finally, we have provided a comprehensive thermodynamic analysis of both extremal and non-extremal RN black holes, which has further interesting implications. For extremal RN black holes, it unexpectedly reveals a connection between super-extremality and the logarithmic entropy corrections, which were previously viewed as two independent research directions. For non-extremal RN black holes, we also demonstrate a relationship between the shift in the charge-to-mass ratio and the thermodynamic stability properties of the black holes. Both of these aspects merit further investigation.
@@ Acknowledgments
We thank Grant Remmen, Nima Arkani-Hamed, and Yu-Tin Huang for informing us of their earlier work on loop effects of massless particles and for the constructive communications after our work was posted on arXiv. XY would like to thank Yu Tian and Hongbao Zhang for useful discussions. YX is also thankful to the Higgs Centre for Theoretical Physics at the University of Edinburgh for providing research facilities and hospitality during the visit. This work was supported in part by China Scholarship Council with Grant No. 202408130101, the Hebei Natural Science Foundation with Grant No. A2024201012, and the National Natural Science Foundation of China with Grant No. 12475048.